Which of the following exponential smoothing constant values puts the same weight on the most recent time series value as does a 3-period moving average?

a. .2
b. .25
c. .75
d. .8
Answer #1

exponential smoothing constant values puts the same weight on the most recent time series value as does a 3-period moving average is

Alpha=.20

Calculation..-as (m+1)/2=3..so from here m=5...lag is 5...so value of constant=1/m=1/5=.2(ans)

Note-if there is any understanding problem regarding this please feel free to ask via comment box..thank you